OpenStudy (thatcreepyboy):

If anyone can help me I'll medal and fan. Find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. Use 3.14 for pi. 3 inches 4 inches 8 inches 16 inches

OpenStudy (thatcreepyboy):

Anyone?

OpenStudy (isaidavila):

The volumen of a cone is 1/3h(πr^2)=V 1/3(5)(3.14*r^2)=83.74 (83.74*3)/(5π)=r^2

OpenStudy (isaidavila):

So now you know the radius, and the diameter is 2r

OpenStudy (thatcreepyboy):

Would that make it 3?

OpenStudy (isaidavila):

I get r≈4 so diameter is 2(4)≈ 8

OpenStudy (thatcreepyboy):

Okay thank you.